what is the value of changeing c if MPC 0.5 Innestal investment is 1000



1
Expert's answer
2021-06-03T12:31:56-0400

if MPC = 0.5, then the value of multiplier is m = 1/(1 - 0.5) = 2.

So, the investment of 1000 will increase GDP by 1000*2 = 2000 and consumption by 2000 - 1000 = 1000.